The group president for Household Products Corporations was reviewing price and promotion alternatives for two products: Green-Mist and Green-Rinse. Both products were designed to keep the bathroom odor free---Green-Mist being an aerosol deodorant and Green-Rinse being an in-tank toilet rinse.

      The price and promotion alternatives recommended for the two products by their respective brand managers included the possibility of using additional promotion or a price reduction to stimulate sales volume. A volume, price, and cost summary for the two products follows:

                  Green-Mist      Green-Rinse

      Unit Price      $5.00      $3.00

      Unit Variable Costs      $3.50      $0.50

      Unit Contribution       $1.50      $2.50

      Unit Volume      1,000,000 units      1,500,000 units

      Both brand managers included a recommendation to either reduce price by 10 percent or invest an incremental $300,000 in advertising

a.       What absolute increase in unit sales and dollar sales will be necessary to recoup the incremental increase in advertising expenditures for Green-Mist? For Green-Rinse?

b.       How many additional sales dollars must be produced to cover each $1.00 of incremental advertising for Green-Mist? For Green-Rinse?

c.       What incremental increase in unit sales and dollar sales will be necessary to maintain the level of total contribution dollars if the price of each product is reduced by 10%?
Answer #1
  1. If $300,000 is invested in advertising, there is a need to make sales of the same amount.

For Green-Mist:

Desired sales (units) = $300,000/$1.5 =200,000 units

Desired sales ($) = 200,000 x $5 =$1,000,000

For Green-Rinse:

Desired sales (units) = $300,000/$2.5 =120,000 units

Desired sales ($) = 120,000 x $3 =$360,000

b)

To cover each $1 of advertising following additional sales amount is needed.

For Green-Mist:

Sales dollar for $300,000 cost = $1,000,000

Additional sales dollars for $1 cost = $1,000,000/300,000 = $3.33

For Green-Rinse:

Sales dollar for $300,000 cost = $360,000

Additional sales dollars for $1 cost = $360,000/300,000 = $1.2

c)

If price is reduced by 10%

For Green-Mist:

New price = $5 – 10% of $5 = $4.5

New contribution = $4.5 - $3.5 = $1

Total contribution required = $1.5 x 1,000,000 = $1,500,000

Units = $1,500,000/$1 = 1,500,000

Incremental units = 500,000 units

Incremental dollar sales = 500,000 x $4.5 = $2,250,000

For Green-Rinse:

New price = $3 – 10% of $3 = $2.7

New contribution = $2.7 - $.5 = $2.2

Total contribution required = $2.5 x 1,500,000 = $3,750,000

Units = $3,750,000/$2.2 = 1,704,545

Incremental units = 204,545 units

Incremental dollar sales = 204,545 x $2.7 = $552,271